OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Dolphin Tricks & Tours: Age Ticket Price Under three years free Three to 12 years $35 Over 12 years $46 On a certain day, the company took 88 people on whale watching trips. There were 44 children aged 12 and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

(44 + x)35 + (88 - (44 + x)) • 46 = C (88 - x)35 + (88 - 44) • 46 = C (44 - x)35 + (88 - 44) • 46 = C (88 + x)35 + (88 - (44 + x)) • 46 = C

OpenStudy (anonymous):

@phi Plz help

OpenStudy (anonymous):

omg i was just doing this question right now! i got the third one.

OpenStudy (anonymous):

was it right

OpenStudy (anonymous):

here this might help :) https://answers.yahoo.com/question/index?qid=20111118131247AAFixjD

OpenStudy (anonymous):

i just need help with 2 more i dont understand

OpenStudy (anonymous):

Solve for x. -ax + 2b > 8 x < the quantity negative 2 times b plus 8 all over negative a x > the quantity negative 2 times b plus 8 all over negative a x < the quantity 2 times b minus 8 all over negative a x > the quantity negative 2 times b minus 8 all over a

OpenStudy (anonymous):

idk im not done yet with the test. but, hey im really goof when it comes to MATH

OpenStudy (anonymous):

^good i meant

OpenStudy (anonymous):

The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F: C = 5 over 9(F – 32) Which of the following formulas correctly solves for F? F = 9 over 5C + 32 F = 9 over 5C – 32 F = 9C + 32 over 5 F = 9C - 32 over 5

OpenStudy (anonymous):

those two im lost on and im not

OpenStudy (anonymous):

for the second question u would have to do the following in order to isolate \[x\]

OpenStudy (anonymous):

here this is how: -ax + 2b > 8 -2b -2b --------------- -ax > 8 - 2b-------> This is what you get after subtracting 2b from both sides

OpenStudy (anonymous):

wouldnt it be 6

OpenStudy (anonymous):

x > the quantity negative 2 times b plus 8 all over negative a

OpenStudy (anonymous):

wait sorry meant this x > the quantity negative 2 times b minus 8 all over a

OpenStudy (anonymous):

-ax > 8 - 2b Then divide by -a to both sides in order to solve x -ax/ -a would leave x but itself. Then divide -a to the right side

OpenStudy (anonymous):

so now u would get this: \[x> (8-2b)/-a\]

OpenStudy (anonymous):

sorry the > would flip over <

OpenStudy (anonymous):

\[x < (8-2b)/-a\]

OpenStudy (anonymous):

the answer to this would be the first option actually. Because when you multiply or divide by a negative number or variable, the inequality is flipped over. In this case, we divided -a, that is why the > sign changed into the < sign.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome :)

OpenStudy (anonymous):

same thing with 46?

OpenStudy (anonymous):

what do you mean by 46? do you mean 4b?

OpenStudy (anonymous):

The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F: C = 5 over 9(F – 32) Which of the following formulas correctly solves for F? F = 9 over 5C + 32 F = 9 over 5C – 32 F = 9C + 32 over 5 F = 9C - 32 over 5

OpenStudy (anonymous):

@phi @cwrw238

OpenStudy (anonymous):

the answer would be like this 9/5(C) + 32 = F